OpenStudy (anonymous):

what is sin^-1(sin (pi/11)

OpenStudy (luigi0210):

Wouldn't it be pi/11? arcsin has a domain of -1 to 1

OpenStudy (anonymous):

thank you, so would tan^-1(tan-pi/3) = -pi/3

OpenStudy (luigi0210):

wait.. is this trig or calculus? :/

OpenStudy (anonymous):

this it trig

OpenStudy (luigi0210):

@katmartins trig or calc?

OpenStudy (luigi0210):

Haha, smartypants :P

OpenStudy (jhannybean):

^___^

OpenStudy (anonymous):

@Luigi0210 this is trig

OpenStudy (anonymous):

it is \(\frac{\pi}{11}\) since \[-\frac{\pi}{2}<\frac{\pi}{11}<\frac{\pi}{2}\]

OpenStudy (luigi0210):

yea, arctans domain is what satellite has shown

OpenStudy (anonymous):

same for tan^-1(tan-pi/3)? does that equal -pi/3

OpenStudy (luigi0210):

Hm, I guess so